OpenStudy (anonymous):

if Z= eˆx (x cos y - y sin y) show that ∂²z/∂x² + ∂²z/∂y² = 0

OpenStudy (anonymous):

Use ln on both sides of the equation to remove the 'e' Now use the derivative implicitly find dx and likewise dy I hope it helped good studies

OpenStudy (anonymous):

∂z/∂x = e^x (xcosy - ysiny) + e^xcosy ∂²z/∂x²=e^x(xcosy -ysiny) + e^x cosy +e^x cosy =xe^x cosy -ye^x siny + 2e^x cosy ∂z/∂y=e^x(-xsiny -siny -ycosy) ∂²z/∂y²= e^x (-xcosy -cosy -cosy +ysiny) = -xe^x cosy -2e^x cosy +ye^x siny ∂²z/∂x² +∂²z/∂y² = (xe^x cosy -ye^x siny + 2e^x cosy) + (-xe^x cosy -2e^x cosy +ye^x siny) = 0 (proved)